1 minus radical3 all divided by 3radical2 minus radical6.

I am trying to rationalize the denominator and am totally lost. Will somebody please send me step by step directions and a final answer please? Thank you!

Answer by lwsshak3(11628)

1 minus radical3 all divided by 3radical2 minus radical6
**
(1-√3)/(3√2-√6)
multiply top and bottom by (3√2+√6)
(1-√3) (3√2+√6)/(3√2-√6) (3√2+√6)
calculating numerator and denominator separately
numerator:
FOIL
(1-√3) (3√2+√6)=3√2+√6-3√6-√18
=3√2-2√6-3√2=-2√6
..
Denominator:
(3√2-√6) (3√2+√6)
Difference of squares
18-6=12
..
(1-√3)/(3√2-√6)=-2√6/12=-√6/2